# Clinicians’ Voice: Fundamental Considerations for XAI in Healthcare

### v1

@@ 4.3 Attitude towards (X)AI in healthcare
However, interviewees also voiced concerns that the distance between them and their patients will become larger when using AI-based DSTs. The same case may have different outcomes because of the patients’ wishes, beliefs, and preferences – “soft features” (interview 2) that are not captured in the data. Personalised treatment is only possible if you interact and connect to your patient, know their story, and incorporate their wishes. Using AI-based devices for some tasks bears the fear that the distance between doctors and their patients becomes so large that they will be dependent on these devices. This outcome would take away a big part of what drives a lot of clinicians, which is simply the interaction with their patients. However, it is important to note that other interviewees, when asked about their opinion on this aspect, said that they are not concerned about it because regardless of where any medical information comes from, it is still them (the clinicians) who are in charge of the situation and in control of the decision. As the name suggests, a DST is merely there to give support – the decision still lies with the clinician who takes them in consultation with the patient.
@@
Bias and AI literacy are other aspects that concern clinicians. In the context of AI, bias relates to the data the model was trained on, and has been an important topic in recent research, especially after criticisms were voiced against some algorithms implemented in practice; see, for example, BBC (2018). Bias in AI exists when the model favors certain people, or provides more accurate predictions for a certain group of individuals. In healthcare, this could mean that an AI-based DST may provide more accurate guidance for a certain group of patients. For example, if the dataset used for training included a higher proportion of old people, the model might give misleading or incorrect results for younger patients. If such systems were to be used, clinicians should be aware of these drawbacks to accurately interpret the output given by the device. Several interviewees pointed this out and referred to it as validity. In fact, five interviewees mentioned that the validation of AI-based devices is extremely important and difficult, making it hard to develop models that can be applied to all patients. Overall, interviewees in this study showed that they are aware of the issue, and further pointed out that this is especially concerning if clinicians rely too much on the AI model instead of using it complementary to their expert knowledge. This concern was voiced especially for younger generations of clinicians who are and will be developing their expertise while these systems are around, potentially making them less critical towards them.
@@
Generally, participants seem to agree that AI is the future and can solve many problems. However, is this really true? To quote one of the interviewees, AI is “kind of the Holy Grail to a lot of people” (interview 6), without considering the context. In a significant number of cases, expert knowledge and/or simple models may be just as good as complex AI models, with the added benefit of providing concrete insights into the decision-making process Semenova et al. (2022), which is also something several interviewees acknowledged.
@@ 4.4 What’s best – feature importances, counterfactual explanations, or rule-based methods?
Comparing only feature importance scores, counterfactual explanations and rule-based models, there is a clear tendency for people to prefer feature importance scores. Interviewees generally agreed that it is intuitive, quick and easy to understand, as you can grasp it all with one glance. It is so easy that “you don’t need to be a professional to understand what’s going on here” (interview 3). While this is clearly a pro of this method, it also raises concerns – “the […] question is, is it really then that simple?” (interview 5). Especially in complex decisions that require consideration of a multitude of factors, it is almost “not believable” (interview 1). Interviewees also pointed out that they are familiar with this method of explaining the output of AI models, and hence may be biased in preferring it over the other two methods.
Counterfactual explanations require a bit more time to be understood, however they can provide one with more in-depth insights. As one interviewee put it, this type of explanation is more “fluid” (interview 5), rather than static as the feature importance scores, and it shows the effect of something changing. This property could also be used to improve prognosis, which however raises the question of causality – a change in features leads to a different prediction by the model, which does not always translate to reality. Lastly, one interviewer explained that looking at counterfactuals means one is essentially looking at other patients, and not focusing on the current patient anymore. On top of that, there are endless other patients out there, so why would one look at this specific counterfactual?
@@
Regardless of the type of explanation, visual explanations are preferred for being easily understood on the spot, and because it’s a way to remove any potential language barriers. Another aspect interviewees raised was uncertainty of the model prediction – being shown how certain the model is about a prediction makes it more trustworthy and helps clinicians navigate the usage of the information given. This ties in with knowing more about the background of the model, such as which data were used for training and testing, how it was validated, how big the sample was, and so on, which came up in about half of the interviews.
@@ 4.5 The role of trust in AI
One third of the interviewees mentioned trust comes automatically when a device is tested, certified, and approved. This is comparable to trusting lab results and equipment without exactly knowing how they work, or when they have last been calibrated. Somewhat in line with that, trust develops over time as you use a device and experience its benefits. Some people are hesitant towards new products, known in economics and sociology as late majority and laggards Rogers (2003). Especially older clinicians seem to be reluctant to use new AI-based medical devices. One interviewee raised the point that trust is distinct from responsibility, that is you can trust a device but in the end you have to take responsibility for your decisions, and “how are you going to feel okay with that responsibility when it’s not your own thought?” (interview 3).
@@ 5 DISCUSSION
5.1 Opportunities for (X)AI research:
Applications in administrative tasks, risk estimation and warning systems, and out-of-hospital health monitoring
: The interviews have shown a positive attitude towards further (X)AI development for clinical practice. Application directions that are mentioned more frequently are lowering the administrative burden for medical personnel, forecasting complications and estimating potential risks and risk prevention, early warning systems, and tools that help monitor patients while they are out-of-hospital, for example via wearable devices.
Development of visual explanations for AI models:
@@ 5.1 Opportunities for (X)AI research:
5.2 Requirements for successful implementation of (X)AI in practice:
More emphasis on multidisciplinary projects is needed for the development of (X)AI techniques:
@@ 5.2 Requirements for successful implementation of (X)AI in practice:
While there is a large body of research on (X)AI-based tools for clinical practice, only very few make it to the bedside. Many reasons for this slow implementation can be attributed to a mismatch between all the different parties involved. Multidisciplinary collaboration is key to building a successful product, and Van de Sande et al. (2022) propose a structured approach to and discuss steps that are needed to develop and safely implement AI in medicine.
(X)AI training is required for clinicians:
@@
Most interviewees in this study have some knowledge about (X)AI, however, mainly for their own interest in the topic. As it stands momentarily, medical education in the Netherlands does not include obligatory classes on AI or AI-based medical devices, which can have significant consequences. First, understanding the background of the model, such as the training data used and how it was validated, is important to judge its performance and get a sense of trust. Second, some explanations may be too simplistic and giving a false sense of understanding, like feature importance plots. It is important for the user to understand what explanations really show and how to interpret them. Third, it is often mistaken that explanations imply causality. Using AI-based tools without sufficient background knowledge makes the users prone to misinterpret the information provided.
@@
Besides, without proper training, we might run the risk of two groups forming among clinicians: those who readily accept and trust AI-based tools, and those who are too resistant to use it. The first group might simply not be aware of important background information that could impact the usage and interpretation of the output. The second group might be too skeptical and hesitant to use technology they do not understand. Both could end up in effects opposite to what we want to achieve with the implementation of AI-based DSTs in clinical practice. Hence, education is key to a successful implementation of these tools as acknowledged in previous work Tjoa and Guan (2021).
(X)AI tools should include more than just explanations:
@@
It is apparent that clinicians desire more than a simple explanation of the prediction, as many have pointed out that they would like to see a certainty score, which is also in line with findings of previous work Tonekaboni et al. (2019). Besides, we can confirm findings of previous work Cai et al. (2019); Lee et al. (2024) that suggest that clinicians are interested in the background of the model, such as information about the training data or benchmark performances. Transparency on training data and representativeness are also important to build trust in the model. A comparison can be made with the setup of clinical trials and evidence-based medicine.
We need more awareness of change management and how to navigate it:
@@
The world is changing at a high pace and especially technological progress is faster than ever, but humans are known to resist change and adapt slowly. From a neurological perspective, humans have been shown to respond faster to familiar stimuli Yang and Beck (2023) and hence tend to stick to droutines. In the workplace, this means that we tend to resist changing our routine or workflow that we are used to. A lot of barriers in the adoption of AI-based DSTs are psychological in nature, as one of the interviewee says (interview 8), and hence we need to draw on knowledge from that field to overcome these barriers. Hence, the development of products for clinical practice needs to be linked to research on human cognition and psychology to allow for a fast and smooth adoption of the technology.
@@
5.3 Warnings regarding (X)AI usage and development:
Black-box models should not be used if they are not a significant improvement over glass-box models:
AI is supposed to be the future, but is it a good idea to use it everywhere? Lately, the field receives more and more criticism for being blindly applied to any project, just because it is AI and it looks good. However, there are many situations where a much simpler model can achieve the same level of performance as a complex black-box model Semenova et al. (2022). Especially when it comes to high stakes situations, one should first use an interpretable glass-box model which allows clinicians to understand the mechanism, and only resort to more complex AI models if they sufficiently outperform glass-box models.
AI should not increase the distance between clinicians and patients:
@@ 6 CONCLUSION
Appendix A METHODOLOGY
We conducted 11 semi-structured interviews with open-ended questions with clinicians trained and working in the Netherlands. In this section we discuss recruitment strategy, interview procedure, material and limitations of our design. The supplementary materials for the interviews are available in the appendix. The study was approved by the ethical body of our institution [to be filled in more detailed after review to preserve this as an anonymous submission]. Interviewees were informed of the goal and procedure of the interviews and signed an informed consent form prior to their participation.
A.1 Participants and recruitment
The main inclusion criterion for participants was that they had successfully completed their medical study in the Netherlands. The sample included clinicians from different fields within medicine (general practice, (pediatric) oncology, radiology, general health psychology, obstetrics and gynaecology, surgery) and with years of experience in their fields varying from two to 30 years. Participants were recruited through networking and word of mouth. None of the interviewees was personally known to the interviewer before the interview was conducted.
A.2 Procedure
Potential interviewees were contacted via email, which stated the reason for reaching out, the person that had provided the contact details, practical information such as the expected time, and the informed consent form. If interviewees responded positively, a follow-up exchange followed to arrange a meeting. Interviews were scheduled for 30 minutes, and the majority of them took place online as participants come from across the Netherlands. Participants were asked to sign the informed consent form prior to starting the interview. During the interview, the interviewer confirmed with participants that they agree to having the interview audio-recorded. Interviewees could ask questions or withdraw from the interview at any moment, and they were explicitly given the opportunity to add any information they wished to mention or ask questions at the end of the interview. All interviews were conducted by the same interviewer to avoid potential differences due to interviewing style.
A.3 Materials
An interview scheme with guiding questions was developed to provide structure and guidance to the interviews (see Table 1). At the start of each interview, the interviewer provided context for the interview after introductions, similar to what participants had read in the informed consent form beforehand. Then, the first set of questions was aimed at inquiring about the interviewee’s background, including the field they specialized in and experience in the field. The following questions were aimed at understanding the current use of AI in their field and healthcare in general. Besides, most interviewees are involved in research that oftentimes includes research on AI for a specific task in their field, and we asked them to share more about that, or other research they’re aware of. Then, we also asked them to come up with use cases of AI in their field which they may find helpful in their job. We also asked what they think are the reasons why these do not yet exist, and why implementation of AI in healthcare in general is slow. The next set of questions was focused on uncovering interviewee’s attitude towards AI in healthcare. This entailed asking specifically for their positive and negative feelings towards using AI in healthcare, and if they could describe their hopes and concerns. In this part of the interview, we also asked them about the role they believe trust plays for the implementation and adoption of (X)AI in practice. Then, for the next set of questions, the interviewer explained three different types of explanations using a simple toy example for illustration (see Appendix B) and asked interviewees to explain their thoughts considering each of these explanations. The explanations that we chose were (1) feature importance scores, (2) counterfactual explanations, and (3) a glass box model. The first method was chosen as feature importance is arguably the most popular method within XAI momentarily. Counterfactual explanations have received considerable attention lately, with many papers proposing algorithms to generate these types of explanations (see, for example, Russell, 2019; Ustun et al., 2019; Kanamori et al., 2021a, b; Karimi et al., 2020, 2021; Mothilal et al., 2020; Mahajan et al., 2019; Maragno et al., 2022). Research from the social and cognitive sciences shows that counterfactual explanations mimic the way humans give explanations themselves, and hence are expected to be good explanations in XAI (Byrne, 2019; Miller, 2019). Lastly, as a contrast, we opted for a glass-box model (weighted rules), such as those generated by Röber et al. (2024), as the third type of explanation. Besides their opinion on these specific explanation types, we included a question on what characteristics (e.g., visual, interactive) they would like to see in an AI system, or what type of information they would like to be shown. Then, at the end of the interview, interviewees were asked if they would like to share any other considerations with us.
A.4 Analysis
Upon the completion of all interviews, each one was transcribed and coded by the interviewer. This involved creating a coding scheme based on the interview scheme and the topics covered in most interviews. The coding scheme was then populated with relevant quotes and excerpts from the transcriptions to allow for a nuanced understanding of the interviewees’ responses. As this was an exploratory study, no predefined coding categories were used; instead, responses were carefully analyzed to identify recurring themes and patterns and the codes were developed organically in response to the data.
A.5 Limitations
The nature of our study bears some limitations we should consider when interpreting the results. First, interviews are known to bear threats to validity and reliability Alshenqeeti (2014). As a response to this, we chose to have the same interviewer conduct all interviews. Interviews were further audio-recorded and then transcribed for the analysis, by that avoiding misinformation through note-taking or incorrect understanding by the interviewer. Furthermore, during the interviews, time was taken to clarify any unclarities to avoid misunderstanding and misinterpretation of the responses. Second, as an explorative interview study, the sample size is limited and we did not perform any quantitative analyses of the results and should be careful with making generalizations. Third, the interviews were scheduled to last 30 minutes, which may not be considered long enough to discuss different aspects of XAI. As this is an exploratory study and clinicians are known to have limited time, we were constrained by this duration. Lastly, we reached out to candidates via email, and in some cases a common connection made the introduction. Generally, it can be expected that individuals interested in XAI are more likely to positively respond to an interview request. Hence, we may only have interviewed individuals with strong (positive or negative) opinions on the topic, by that missing out on the broader group of clinicians.
Appendix B INTERVIEW ADDITIONAL MATERIAL

### v2

@@ 4.3 Attitude towards (X)AI in healthcare
However, interviewees also voiced concerns that the distance between them and their patients will become larger when using AI-based DSTs. The same case may have different outcomes because of the patients’ wishes, beliefs, and preferences – “soft features” (interview 2) that are not captured in the data. Personalized treatment is only possible if you interact and connect with your patient, know their story, and incorporate their wishes. Using AI-based devices for some tasks bears the fear that the distance between doctors and their patients becomes so large that they will be dependent on these devices. This outcome would take away a large part of what drives many clinicians, which is simply the interaction with their patients. However, it is important to note that other interviewees, when asked about their opinion on this aspect, said that they are not concerned about it because regardless of where any medical information comes from, it is still them (the clinicians) who are in charge of the situation and in control of the decision. As the name suggests, a DST is merely there to give support – the decision still lies with the clinician who takes them in consultation with the patient.
@@
Bias and AI literacy are other aspects that concern clinicians. In the context of AI, bias relates to the data the model was trained on, and has been an important topic in recent research, especially after criticisms were voiced against some algorithms implemented in practice; see, for example, BBC (2018). Bias in AI exists when the model favors certain people, that is an AI-based DST may provide more accurate guidance for a certain group of patients. For example, if the dataset used for training included a higher proportion of elderly people, the model might give misleading or incorrect results for younger patients. If such systems were to be used, clinicians should be aware of these drawbacks to accurately interpret the output given by the device. Several interviewees pointed this out and referred to it as validity. In fact, five interviewees mentioned that the validation of AI-based devices is extremely important and difficult, making it hard to develop models that can be applied to all patients. Overall, interviewees in this study showed that they are aware of the issue, and further pointed out that this is especially concerning if clinicians rely too much on the AI model instead of using it complementary to their expert knowledge. This concern was voiced especially for younger generations of clinicians who are and will be developing their expertise while these systems are around, potentially making them less critical towards them.
@@
Generally, participants seem to agree that AI is the future and can solve many problems. However, is this really true? To quote one of the interviewees, AI is “kind of the Holy Grail to a lot of people” (interview 6), without considering the context. In a significant number of cases, expert knowledge and/or simple models may be just as good as complex AI models, with the added benefit of providing concrete insights into the decision-making process (Semenova et al., 2022), which is also something several interviewees acknowledged.
@@ 4.4 What’s best – feature importances, counterfactual explanations, or rule-based methods?
Comparing only feature importance scores, counterfactual explanations and rule-based models, there is a clear tendency for people to prefer feature importance scores. Interviewees generally agreed that they are intuitive, quick and easy to understand, as you can grasp them all with one glance. It is so easy that “you don’t need to be a professional to understand what’s going on here” (interview 3). While this is clearly a pro of this method, it also raises concerns – “the […] question is, is it really then that simple?” (interview 5). Especially in complex decisions that require consideration of a multitude of factors, it is almost “not believable” (interview 1). Interviewees also pointed out that they are familiar with this method of explaining the output of AI models, and hence may be biased in preferring it over the other two methods.
Counterfactual explanations require a bit more time to be understood, however they can provide one with more in-depth insights. As one interviewee put it, this type of explanation is more “fluid” (interview 5), rather than static as the feature importance scores, and it shows the effect of something changing. This property could also be used to improve prognosis, which however raises the question of causality – a change in features leads to a different prediction by the model, which does not always translate to reality. Lastly, one interviewer explained that looking at counterfactuals means one is essentially looking at other patients, and not focusing on the current patient anymore. On top of that, there are endless other patients out there, so why would one look at this specific counterfactual? This statement can be linked to discussions of distance measures in technical papers on counterfactual explanations. That is, the choice of distance function influences which counterfactual is found (Guidotti, 2024). Besides, most approaches use simple distance functions (e.g., Maragno et al., 2022, 2024; Wachter et al., 2017; Kanamori et al., 2020), however ‘closeness’ in real situations may differ from mathematical definitions.
@@
Regardless of the type of explanation, visual explanations are preferred for being easily understood on the spot, and because it is a way to remove any potential language barriers. Another aspect interviewees raised was uncertainty of the model prediction – being shown how certain the model is about a prediction makes it more trustworthy and helps clinicians navigate the usage of the information given. This ties in with knowing more about the background of the model, such as which data were used for training and testing, how it was validated, how big the sample was, and so on, which came up in about half of the interviews.
@@ 4.5 The role of trust in AI
One third of the interviewees mentioned trust comes automatically when a device is tested, certified, and approved. This is comparable to trusting lab results and equipment without exactly knowing how they work, or when they have last been calibrated. Somewhat in line with that, trust develops over time as you use a device and experience its benefits. Some people are hesitant towards new products, known in economics and sociology as late majority and laggards (Rogers, 2003). Especially more senior clinicians seem to be reluctant to use new AI-based medical devices. One interviewee raised the point that trust is distinct from responsibility; that is, you can trust a device but in the end you have to take responsibility for your decisions, and “how are you going to feel okay with that responsibility when it’s not your own thought?” (interview 3). This adds a new angle to the discussion on trust of AI models that has not yet been sufficiently addressed in research.
@@ 5 Discussion
5.1 Opportunities for (X)AI research
Applications in administrative tasks, risk estimation and warning systems, and out-of-hospital health monitoring.
The interviews have shown a positive attitude towards further (X)AI development for clinical practice. Application directions that are mentioned more frequently are lowering the administrative burden for medical personnel, forecasting complications and estimating potential risks and risk prevention, early warning systems, and tools that help monitor patients while they are out-of-hospital, for example via wearable devices.
Development of visual explanations for AI models.
@@ 5.1 Opportunities for (X)AI research
5.2 Requirements for successful implementation of (X)AI in practice
More emphasis on multidisciplinary projects is needed for the development of (X)AI techniques.
@@ 5.2 Requirements for successful implementation of (X)AI in practice
While there is a large body of research on (X)AI-based tools for clinical practice, only very few make it to the bedside. Many reasons for this slow implementation can be attributed to a mismatch between all the different parties involved. Multidisciplinary collaboration is key to building a successful product, and Van de Sande et al. (2022) propose a structured approach to develop and safely implement AI in medicine.
(X)AI training is required for clinicians.
@@
Most interviewees in this study have some knowledge about (X)AI, however, mainly for their own interest in the topic. As it stands momentarily, medical education in the Netherlands does not include obligatory classes on AI or AI-based medical devices, which can have significant consequences. First, understanding the background of the model, such as the training data used and how it was validated, is important to judge its performance and get a sense of trust. Second, it is often mistaken that explanations imply causality. Using AI-based tools without sufficient background knowledge makes the users prone to misinterpret the information provided. Third, some explanations may be too simplistic and may give a false sense of understanding, like feature importance plots. Instead, explanations should stimulate critical thinking so that the user can evaluate the given output themselves. Providing too simplistic explanations may hide crucial pieces of information that otherwise would raise doubts. Especially when the provided information is in line with the user’s expectations, simple explanations may increase the risk of confirmation bias. Confirmation bias has indeed been shown to exist in human-technology interaction (e.g., Wan et al., 2022; Bauer et al., 2023) and has been acknowledged in the XAI community as well (e.g., Ghassemi et al., 2021; Rudin, 2019; Cina et al., 2023). On the other hand, experienced users may disregard a simple explanation altogether as they may find it unbelievable because the true reason for a given output is likely not as simple as a feature importance plot might suggest. After all, why are we using AI if it really were this easy?
@@
Besides, without proper training, we might run the risk of two groups forming among clinicians: those who readily accept and trust AI-based tools, and those who are too resistant to use it. The first group might simply not be aware of important background information that could impact the usage and interpretation of the output. The second group might be too skeptical and hesitant to use technology they do not understand. Both could end up in effects opposite to what we want to achieve with the implementation of AI-based DSTs in clinical practice. Hence, training is key to a successful implementation of these tools as acknowledged in the previous work of Tjoa and Guan (2021).
(X)AI tools should include more than just explanations.
@@
It is apparent that clinicians desire more than a simple explanation of the prediction, as many have pointed out that they would like to see a certainty score, which is also in line with observations of Tonekaboni et al. (2019). Additionally, we can confirm findings of Cai et al. (2019); Lee et al. (2024) that suggest that clinicians are interested in the background of the model, such as information about the training data or benchmark performances. Transparency on training data and representativeness are also important to build trust in the model. A comparison can be made with the setup of clinical trials and evidence-based medicine.
We need more awareness of change management and how to navigate it.
@@
The world is changing at a high pace and especially technological progress is faster than ever, but humans are known to resist change and adapt slowly. From a neurological perspective, humans have been shown to respond faster to familiar stimuli (Yang and Beck, 2023), and hence, tend to stick to routines. In the workplace, this means that we tend to resist changing our routine or workflow that we are used to. Many barriers in the adoption of AI-based DSTs are psychological in nature, as one of the interviewee says (interview 8), and hence we need to draw on knowledge from that field to overcome these barriers. Hence, the development of products for clinical practice needs to be linked to research on human cognition and psychology to allow for a fast and smooth adoption of the technology.
@@
5.3 Warnings regarding (X)AI usage and development
Black-box models should not be used if they are not a significant improvement over glass-box models.
AI is supposed to be the future, but is it a good idea to use it everywhere? Lately, the field receives more and more criticism for being blindly applied to any project, just because it is AI and it looks good. However, there are many situations where a much simpler model can achieve the same level of performance as a complex black-box model (Semenova et al., 2022). Especially when it comes to high stakes situations, one should first use an interpretable glass-box model which allows clinicians to understand the mechanism, and only resort to more complex AI models if they sufficiently outperform glass-box models.
AI should not increase the distance between clinicians and patients.
@@ 6 Conclusion
Appendix A Additional interview material